### v3

@@ Abstract
Visualizations play a pivotal role in daily communication in an increasingly data-driven world. Research on multimodal large language models (MLLMs) for automated chart understanding has accelerated massively, with steady improvements on standard benchmarks. However, for MLLMs to be reliable, they must be robust to misleading visualizations, charts that distort the underlying data, leading readers to draw inaccurate conclusions that may support disinformation. Here, we uncover an important vulnerability: MLLM question-answering accuracy on misleading visualizations drops on average to the level of a random baseline. To address this, we introduce the first inference-time methods to improve performance on misleading visualizations, without compromising accuracy on non-misleading ones. The most effective method extracts the underlying data table and uses a text-only LLM to answer the question based on the table. Our findings expose a critical blind spot in current research and establish benchmark results to guide future efforts in reliable MLLMs. [^1] Code and data available at: github.com/UKPLab/arxiv2025-misleading-visualizations
keywords:
large language models, chart understanding, visualization
@@
In an increasingly data-driven world, visualizations are widely used by scientists, journalists, governments, or companies to efficiently communicate data insights to a broad audience [1]. They play a crucial role in crisis settings, such as during the COVID-19 pandemic, where charts were shared daily to inform the public about the spread of the disease [2]. In many cases, visualizations support a message more convincingly than if the underlying data table was shown directly to readers [3]. However, visualizations can also be misleading [4]. This occurs when design flaws, or misleaders, distort the underlying data in a way that prevents a correct interpretation by the reader [5, 6, 4, 7, 8, 9, 10, 11, 12]. These misleaders appear across a wide range of visualizations, including bar, pie, and line charts, choropleth maps, or scatter plots. Recent taxonomies have documented over 70 distinct types of misleaders observed in real-world examples [7, 9]. Major categories include axes manipulations, such as truncated and inverted axes, or the use of visual manipulations like area encoding and 3D effects. Figure 1 shows three real-world examples of misleading visualizations [7] paired with multiple-choice questions. Given a question about the underlying data table, their misleaders lead readers to infer a misleading answer (in purple) instead of the correct one (in green). Figure 2 illustrates how two visualizations of the same data, one non-misleading and the other misleading, can lead to different interpretations by readers.
@@
Figure 1: Three examples of real-world misleading visualizations [7] with multiple choice questions. The correct answer is colored in green, while the wrong answer supported by the misleader is colored in purple.
@@
Figure 2: A pair of bar charts, one non-misleading and one misleading, representing the same underlying data [11]. Both are paired with the same Likert-scale question, where 1 means “a little” and 6 means “a lot”. A consistent interpretation requires identical responses for both visualizations. Assuming a score of 3 is given for the non-misleading chart, only a score of 3 is valid for the misleading one. However, the truncated axis deceives readers in choosing a value higher than 3.
@@
Prior work has revealed the potential of misleading visualizations to deceive human readers. Some studies [6, 13, 14, 15] have shown that such visualizations reduce human accuracy when answering multiple-choice, free-text, or rank questions about the underlying data. Other studies [6, 16, 11] have demonstrated that human readers tend to provide different Likert-scale answers to a question depending on whether they view a misleading or non-misleading visualization of the same data table.
@@
Misleading visualizations pose a serious threat to society. Due to their high deceptive potential, they can be exploited by malicious actors to promote online disinformation [4]. For instance, the visualizations in Figure 1 mislead readers on sensitive topics such as COVID-19, abortion, and gun violence statistics. Misleading visualizations can shift public opinion, even on highly polarized political debates like Brexit [17].
@@
Multimodal large language models (MLLMs) have sparked a massive interest in automated chart understanding research [1], with steady progress on reference benchmarks such as ChartQA [18, 19], highlighting their strong potential as interactive assistants for extracting data insights. However, existing research has largely overlooked the important threat posed by misleading visualizations. If MLLMs, like humans, are easily deceived by misleading visualizations, malicious actors could exploit this vulnerability. MLLMs confronted to misleading visualizations could present incorrect interpretations of the underlying data to users, contributing to the spread of disinformation, and reinforcing human belief in it. These risks underscore the urgency of thoroughly assessing the vulnerability of MLLMs to misleading visualizations and developing effective mitigation strategies.
In this work, we conduct the first comprehensive study to assess and mitigate the vulnerability of 18 MLLMs of varying sizes to misleading visualizations. This collection includes 11 open-weight MLLMs from the Llava-Next [20], Qwen2VL [21], Ovis-1.6 [22], and InternVL2.5 [23] families, ranging from 2 to 38 billion parameters. Additionally, we include four commercial models, GPT4 [24], GPT4o [24], Gemini-1.5-flash [25], and Gemini-1.5-pro [25], as well as three open-weight MLLMs specifically trained for visualization understanding: Llava-Chart-Instruct [26], TinyChart [27], and ChartGemma [28] ranging from 3 to 13 billion parameters.
In Experiment 1, we compare the question-answering (QA) accuracy of MLLMs across three datasets: (a) a misleading visualization dataset containing $n$ = 143 instances, and featuring 17 types of misleaders, defined in Table 1 [13, 14, 15, 7]. It combines the existing synthetic CALVI ($n$ = 45) [13] and CHARTOM ($n$ = 56) [14, 15] datasets with 42 real-world examples [7] that we manually annotated with multiple choice questions; (b) a non-misleading visualization dataset ($n$ = 124) [13, 14, 15, 29], also combining real-world and synthetic cases; and (c) ChartQA [18], the reference real-world non-misleading benchmark ($n$ = 2500). In Experiment 2, we evaluate the consistency of MLLMs, i.e., whether they provide the same response to a Likert-scale question when shown either a misleading or a non-misleading visualization of the same underlying data. We use the synthetic dataset of Lauer & O’Brien, which consists of four such visualization pairs [11]. Detailed statistics about the datasets, including the number of misleader categories and the type of questions, are reported in Table 2.
Experiments 1 and 2 demonstrate that MLLMs, like humans, are indeed vulnerable to misleading visualizations. Experiment 1 yields several key insights: QA accuracy drops by up to 65.5 percentage points (pp) compared to ChartQA, with half the MLLMs failing to outperform the random baseline; MLLMs are less vulnerable to real-world misleading visualizations than to synthetic ones; and vulnerability varies substantially depending on the type of misleader.In addition, Experiment 2 shows that MLLMs exhibit inconsistency when assigning Likert-scale scores to two visualizations of the same data when one of them has a truncated axis.
To mitigate the high vulnerability of MLLMs observed in Experiment 1, we introduce six inference-time correction methods in Experiment 3. These methods modify the input image and/or prompt without requiring additional training data or fine-tuning. We evaluate them on both the misleading and non-misleading visualization datasets used in Experiment 1. The goal is to improve QA accuracy on misleading visualizations while maintaining performance on non-misleading ones.
We compare the following methods, illustrated in Figure 3: (1) including a warning message to the prompt to make the MLLM aware of specific misleaders in the visualization; (2) extracting the axes using the MLLM and incorporating them into the prompt; (3) extracting the underlying data table and adding it into the prompt; (4) combining both extracted axes and table into the prompt; (5) providing the extracted table, without the visualization, to a text-only LLM, reframing the task as table-based QA; and (6) using a text-only LLM to generate code for a new visualization based on the extracted table, which then replaces the original visualization in the QA prompt.
@@
In summary, our contributions are twofold: (1) We present the first extensive analysis of the vulnerabilities of 18 MLLMs to 17 types of misleading visualizations, evaluating both QA accuracy and Likert-scale consistency. (2) We are the first to introduce six correction methods to mitigate the negative impact of misleaders on QA performance, while maintaining accuracy on non-misleading visualizations.
The results of the three experiments expose a critical blind spot in current automated chart understanding research, highlighting the need for future work to make MLLMs more robust to misleading visualizations.
@@
Figure 3: Illustration of the six inference-time correction methods applied on a misleading visualization from CALVI [13]. The visualization suffers from inconsistent tick intervals on the y-axis. The correct answer is True. In this case, it is only predicted after redrawing the visualization.
@@ 1.1 Experiment 1 - Assessing MLLM vulnerabilities (Accuracy)
Figure 4: Top: Accuracy (%) of various MLLMs on misleading visualization [13, 14, 15, 7], non-misleading visualization [13, 14, 15, 29], and ChartQA datasets [18]. The horizontal dashed line indicates the accuracy of the random baseline on misleading visualizations. Models are sorted by increasing accuracy on ChartQA. Bottom: Accuracy (%) of various MLLMs on subsets of the misleading visualizations. The horizontal dashed lines indicate average human accuracy on CALVI and CHARTOM [13, 14, 15].
1.1.1 Performance is worse on misleading visualizations
All MLLMs perform substantially worse on misleading visualization compared to the non-misleading visualizations dataset, with accuracy dropping by up to 34.8 pp, with an average of 23.4 pp. The decline is even more pronounced when compared to ChartQA, reaching up to 65.5 pp, with an average of 54.0 pp. Furthermore, the mean MLLM accuracy on misleading visualizations (25.6%) is on par with the random baseline score (25.6%). These results demonstrate that MLLMs struggle to correctly interpret the distorted data, making them highly vulnerable to misleaders.
For comparison, average human accuracy reported in prior work on CALVI and CHARTOM is 80 and 89% on non-misleading visualizations, dropping to 39 and 49% on misleading ones, respectively [13, 14, 15]. On the same datasets, MLLMs achieve an average of 51 and 47% on non-misleading visualizations and 24 and 21% on misleading ones, falling short of human performance in all cases.
@@ 1.1.1 Performance is worse on misleading visualizations
1.1.2 Performance is higher on real-world misleading visualizations
MLLMs perform better on the real-world subset of the misleading visualizations dataset, achieving an average accuracy of 33.5%, compared to 24% on CALVI and 21% on CHARTOM. In general, the best-performing MLLMs on misleading visualizations, GPT4o, Gemini-1.5-pro, and InternVL2.5-38B, primarily distinguish themselves through their higher performance on this subset. We hypothesize that this advantage stems from their parametric knowledge, which extends beyond 2022, the endpoint of the real-world subset. This allows these models to answer some questions based on stored knowledge about past events and statistics, without relying on the content of the visualization. In such cases, parametric knowledge effectively serves as a form of protection against real-world misleaders. However, this finding should be interpreted with caution. In practical scenarios, the underlying data will often be recent and therefore unlikely to be covered by the model’s parametric knowledge.
1.2 Performance varies by type of misleader
Figure 5: Average Likert-scale ratings (1 to 6) of MLLMs and human readers for the four pairs of misleading and non-misleading visualizations depicting the same underlying data from the dataset of Lauer & O’Brien [11]. MLLM results are reported with standard deviations. Human results are reported from prior work [11].
Not all misleading visualizations are equally deceptive. This was already observed with human readers [13, 14, 15], and is here further confirmed with MLLMs. Detailed accuracy scores per misleader and MLLM are reported in Figure 7. The three misleaders on which MLLMs perform best are concealed uncertainty, misleading annotations, and misrepresentation, with average accuracies of 64.8, 50.0, and 43.7, respectively. The four on which they perform worse are overplotting and inappropriate use of pie charts, both at 0.0%, and area encoding and missing data at 5.6%. For the two most represented misleaders, truncated axis and inverted axis, the MLLMs achieve average accuracies of 23.5 and 27.4 %, respectively.
There are notable differences in which misleaders pose greater challenges to MLLMs compared to humans, and vice versa. For instance, area encoding is highly misleading for MLLMs, but not for human readers, who achieve an average accuracy of 91.5% on this type of misleader [14]. Conversely, both humans and MLLMs struggle with misleaders such as overplotting, missing data, and 3D effects [13]. The use of a dual axis is a misleader that poses greater difficulty for humans, who score an average accuracy of 16.1% on CHARTOM [14], compared to MLLMs, which achieve an average of 35.4%. These results highlight a notable difference in how humans and MLLMs process visualizations and are deceived by misleaders. This suggests that the methods needed to make MLLMs more robust to misleading visualizations will, at least in part, differ from those designed to protect human readers [30].

### v4

@@ Abstract
Visualizations play a pivotal role in daily communication in an increasingly data-driven world. Research on multimodal large language models (MLLMs) for automated chart understanding has accelerated massively, with steady improvements on standard benchmarks. However, for MLLMs to be reliable, they must be robust to misleading visualizations, i.e., charts that distort the underlying data, leading readers to draw inaccurate conclusions that may support disinformation. Here, we uncover an important vulnerability: MLLM question-answering (QA) accuracy on misleading visualizations drops on average to the level of the random baseline. To address this, we introduce the first inference-time methods to improve QA performance on misleading visualizations, without compromising accuracy on non-misleading ones. We find that two methods, table-based QA and redrawing the visualization, are effective, with improvements of up to 19.6 percentage points. We make our code and data available.111github.com/UKPLab/arxiv2025-misleading-visualizations
1 Introduction
@@
In an increasingly data-driven world, visualizations are widely used by scientists, journalists, governments, or companies to efficiently communicate data insights to a broad audience (Huang et al., 2025). They play a crucial role in crisis settings, such as during the COVID-19 pandemic, where charts were shared daily to inform the public about the spread of the virus (Zhang et al., 2021). In many cases, visualizations support a message more convincingly than if the underlying data table was shown directly to readers (Pandey et al., 2014). However, visualizations can also be misleading (Correll and Heer, 2017). This occurs when design flaws, or misleaders, distort the underlying data in a way that prevents a correct interpretation by the reader (Tufte and Graves-Morris, 1983; Pandey et al., 2015; Correll and Heer, 2017; Lauer and O’Brien, 2020; McNutt et al., 2020; Yang et al., 2021; Lo et al., 2022; Lisnic et al., 2023; Lan and Liu, 2025). These misleaders appear across a wide range of visualizations, including bar, pie, and line charts, choropleth maps, or scatter plots. Recent taxonomies have documented over 70 distinct types of misleaders observed in real-world examples (Lo et al., 2022; Lan and Liu, 2025). Major categories include axes manipulations, such as truncated and inverted axes, or the use of visual manipulations like area encoding and 3D effects. Figure 1 shows three real-world examples of misleading visualizations (Lo et al., 2022) paired with multiple-choice questions. Given a question about the underlying data table, their misleaders lead readers to infer a misleading answer instead of the correct one. Figure 2 illustrates how two visualizations of the same underlying data, one non-misleading and the other misleading, can lead to different interpretations by readers.
@@
Figure 1: Three examples of real-world misleading visualizations (Lo et al., 2022) with multiple choice questions. The correct answer is colored in green, while the wrong answer supported by the misleader is colored in purple.
@@
Figure 2: Non-misleading and misleading visualizations of the same data (Lauer and O’Brien, 2020) with a Likert-scale question where 1 means “a little” and 6 means “a lot”. A consistent interpretation requires identical responses for both visualizations. However, the truncated axis deceives readers in choosing a higher value for the misleading visualization.
@@
Prior work has revealed the potential of misleading visualizations to deceive human readers. Some studies (Pandey et al., 2015; Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024) have shown that such visualizations reduce human accuracy when answering multiple-choice, free-text, or rank questions about the underlying data. Other studies (Pandey et al., 2015; O’Brien and Lauer, 2018; Lauer and O’Brien, 2020) have demonstrated that human readers tend to provide different Likert-scale answers to a question depending on whether they view a misleading or non-misleading visualization of the same data table.
@@
Misleading visualizations pose a serious threat to society. Due to their deceptive potential, they can be exploited by malicious actors to promote online disinformation (Correll and Heer, 2017). For instance, the charts in Figure 1 mislead readers on sensitive topics such as COVID-19, abortion, and gun violence. Misleading visualizations can shift public opinion, even on polarized debates like Brexit (Tartaglione and de Wit, 2025).
@@
Multimodal large language models (MLLMs) have sparked a massive interest in automated chart understanding research (Huang et al., 2025), with steady progress on reference benchmarks such as ChartQA (Masry et al., 2022; Hoque et al., 2022), highlighting their strong potential as interactive assistants for extracting data insights. However, existing research has largely overlooked the important threat posed by misleading visualizations. If MLLMs, like humans, are easily deceived by misleading visualizations, malicious actors could exploit this vulnerability. MLLMs confronted to misleading visualizations could present incorrect interpretations of the underlying data to users, contributing to the spread of disinformation, and reinforcing human belief in it.
In this work, we conduct the first comprehensive study to assess and mitigate the vulnerability of 18 MLLMs of varying sizes to misleading visualizations. In Experiment 1, we compare the question-answering (QA) accuracy of MLLMs on misleading and non-misleading visualizations. In Experiment 2, we evaluate the consistency of MLLMs, i.e., whether they provide the same response to a Likert-scale question when shown either a misleading or a non-misleading visualization of the same underlying data. The results of both experiments demonstrate that MLLMs, like humans, are indeed vulnerable to misleading visualizations. In Experiment 3, we introduce six inference-time correction methods to reduce the vulnerability of MLLMs to misleading visualizations in a QA setting. We evaluate them on the datasets of Experiment 1. The goal is to improve QA accuracy on misleading visualizations while maintaining performance on non-misleading ones.
@@
In summary, our contributions are twofold: (1) We present the first extensive analysis of the vulnerabilities of 18 MLLMs to 17 types of misleading visualizations, evaluating both QA accuracy and Likert-scale consistency. (2) We are the first to introduce six correction methods to mitigate the negative impact of misleaders on QA tasks.
2 Methodology
2.1 Datasets
Experiments 1 and 3 We compare the question-answering (QA) accuracy of MLLMs across three datasets: (a) a misleading visualization dataset containing $n$ = 143 instances, and featuring 17 types of misleaders, defined in Table 2 (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024; Lo et al., 2022); (b) a non-misleading visualization dataset ($n$ = 124) (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024; Lee et al., 2017), also combining real-world and synthetic cases; and (c) ChartQA (Masry et al., 2022), the reference real-world non-misleading benchmark ($n$ = 2500). Datasets (a) and (b) combine three existing resources and one introduced in this work. First, CALVI (Ge et al., 2023) includes 45 misleading and 15 non-misleading visualizations based on synthetic data, each paired with a multiple-choice question (MCQ). The dataset was initially designed to evaluate humans. Second, CHARTOM (Rho et al., 2023; Bharti et al., 2024) contains 56 samples, including 28 MCQs, 20 free-text questions, and 8 rank questions. Each question is linked to two visualizations, one misleading and one non-misleading. Like CALVI, the underlying data is synthetic, and the test was originally designed to evaluate humans. Third, VLAT (Lee et al., 2017), the reference dataset to evaluate humans, provides 12 non-misleading visualizations, each paired with three to seven MCQs, for a total of 53 instances. The visualizations are based on real-world data. Fourth, we introduce a dataset of 42 real-world misleading visualizations, each annotated with a MCQ with three to four choices. They come from a collection annotated with the misleaders affecting them (Lo et al., 2022), which inspired the synthetic examples in CALVI. We manually create MCQs, using as reference those from CALVI and CHARTOM. By incorporating visualizations and questions about real-world data, we introduce direct conflicts with MLLMs’ parametric knowledge, allowing us to assess their vulnerability in real-world scenarios.
Experiment 2 We use four pairs of visualizations, one misleading and the other non-misleading, originally designed by Lauer and O’Brien (2020) to evaluate humans. The pairs cover three distinct misleaders.
Table 1 provides detailed dataset statistics.
Dataset Experiments $n$ $m$ Question types
Misleading visualizations 1 & 3 143 17 MCQ, Free-text, rank
$\hookrightarrow$ CALVI (misleading) 1 & 3 45 14 MCQ
$\hookrightarrow$ CHARTOM (misleading) 1 & 3 56 7 MCQ, Free-text, rank
$\hookrightarrow$ Real-world 1 & 3 42 12 MCQ
Non-misleading visualizations 1 & 3 124 - MCQ, Free-text, rank
$\hookrightarrow$ CALVI (non-misleading) 1 & 3 15 - MCQ
$\hookrightarrow$ CHARTOM (non-misleading) 1 & 3 56 - MCQ, Free-text, rank
$\hookrightarrow$ VLAT 1 & 3 53 - MCQ
ChartQA 1 2500 - Free-text
Lauer & O’Brien 2 8 3 Likert-scale
Table 1: Datasets statistics. $n$ is the number of instances. $m$ is the number of misleader categories.
2.2 Correction methods
@@
Figure 3: Illustration of the six inference-time correction methods applied to a misleading visualization from CALVI (Ge et al., 2023). The visualization suffers from inconsistent tick intervals on the y-axis.
In Experiment 3, we compare six inference-time correction methods, illustrated in Figure 3, to increase QA accuracy on misleading visualization without compromising it on non-misleading ones.
(1) Misleader warning: we insert in the prompt a message to make the MLLM aware of the misleader in the visualization. The message is identical for all instances with the same misleader.
Providing (2) the axes data, (3) the data table, (4) or both: we prompt the MLLM in zero-shot to extract the axes labels or underlying data table. The axes and tables are formatted as text strings. We do not impose constraints on the delimiters used to indicate new rows and columns. The axes, the table, or both are then provided as additional input in the prompt.
(5) Table-based QA: after extracting the table with the MLLM, we provide it alone to a text-only LLM, reframing the task as table-based QA.
(6) Redrawn visualization: after extracting the table with the MLLM, we provide it to a text-only LLM which generates Python code to create a visualization using Matplotlib (Hunter, 2007). If the code compiles successfully, the generated visualization replaces the original one in the QA prompt; otherwise, the original one is used.
2.3 Implementation details
We conduct experiments with 18 MLLMs, including 11 open-weight MLLMs from the Llava-Next (Liu et al., 2024), Qwen2VL (Wang et al., 2024), Ovis-1.6 (Lu et al., 2024), and InternVL2.5 (Chen et al., 2024) families, ranging from 2 to 38 billion parameters. Additionally, we include four commercial models, GPT4 (OpenAI, 2023), GPT4o (OpenAI, 2023), Gemini-1.5-flash (Gemini-Team, 2024), and Gemini-1.5-pro (Gemini-Team, 2024), as well as three open-weight MLLMs specifically trained for visualization understanding: Llava-Chart-Instruct (Zeng et al., 2025), TinyChart (Zhang et al., 2024), and ChartGemma (Masry et al., 2025) ranging from 3 to 13 billion parameters. Qwen2.5-7B (Qwen-Team, 2024) serves as the text-only LLM. We use the transformers library (Wolf et al., 2020) to access all open models and the Azure API and Google Cloud API for GPT and Gemini models, respectively. We set the random seed at 42 and the temperature to 0. Following the standard ChartQA setup, all (M)LLMs are prompted in zero-shot. We report TinyChart results with the Direct approach (Zhang et al., 2024).
@@ 3.1 Experiment 1 - Assessing MLLM vulnerabilities (Accuracy)
Figure 4: Top: Accuracy (%) of various MLLMs on misleading visualization, non-misleading visualization, and ChartQA datasets. The horizontal dashed line indicates the accuracy of the random baseline on misleading visualizations. Models are sorted by increasing accuracy on ChartQA. Bottom: Accuracy (%) of various MLLMs on subsets of the misleading visualizations. The horizontal dashed lines indicate average human accuracy on CALVI and CHARTOM (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024).
@@
Performance is worse on misleading visualizations. All MLLMs perform substantially worse on misleading visualization compared to the non-misleading visualizations dataset, with accuracy dropping by up to 34.8 percentage points (pp), with an average of 23.4 pp. The decline is even more pronounced compared to ChartQA, reaching up to 65.5 pp, with an average of 54.0 pp. Furthermore, the mean MLLM accuracy on misleading visualizations (25.6%) is on par with the random baseline score (25.6%). These results demonstrate that MLLMs struggle to correctly interpret data distorded by misleaders.
For comparison, average human accuracy reported on CALVI and CHARTOM is 80 and 89% on non-misleading visualizations, dropping to 39 and 49% on misleading ones, respectively (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024). On the same datasets, MLLMs achieve an average of 51 and 47% on non-misleading visualizations and 24 and 21% on misleading ones, falling short of human performance in all cases.
@@
Performance is higher on real-world misleading visualizations than on other subsets. MLLMs perform better on the real-world subset of the misleading visualizations dataset, achieving an average accuracy of 33.5%, compared to 24% on CALVI and 21% on CHARTOM. In general, the best-performing MLLMs on misleading visualizations, GPT4o, Gemini-1.5-pro, and InternVL2.5-38B, primarily distinguish themselves through their higher performance on this subset. We hypothesize that this advantage stems from their parametric knowledge, which extends beyond 2022, the endpoint of the real-world subset. This allows these models to answer some questions based on stored knowledge about past events and statistics, without relying on the content of the visualization. In such cases, parametric knowledge effectively serves as a form of protection against real-world misleaders. However, this finding should be interpreted with caution. In practical scenarios, the underlying data will often be recent and therefore unlikely to be covered by the model’s parametric knowledge.
Figure 5: QA Accuracy (%) per type of misleader.
Performance varies by type of misleader. Not all misleading visualizations are equally deceptive. This was already observed with human readers (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024), and is here further confirmed with MLLMs. Detailed accuracy scores per misleader and MLLM are reported in Figure 5. The three misleaders on which MLLMs perform best are concealed uncertainty, misleading annotations, and misrepresentation, with average accuracies of 64.8, 50.0, and 43.7, respectively. The four on which they perform worse are overplotting and inappropriate use of pie charts, both at 0.0%, and area encoding and missing data at 5.6%. For the two most represented misleaders, truncated axis and inverted axis, the MLLMs achieve average accuracies of 23.5 and 27.4 %, respectively. There are notable differences in which misleaders pose greater challenges to MLLMs compared to humans, and vice versa. For instance, area encoding is highly misleading for MLLMs, but not for human readers, who achieve an average accuracy of 91.5% on this type of misleader (Rho et al., 2023). Conversely, both humans and MLLMs struggle with misleaders such as overplotting, missing data, and 3D effects (Ge et al., 2023). The use of a dual axis is a misleader that poses greater difficulty for humans, who score an average accuracy of 16.1% on CHARTOM (Rho et al., 2023), compared to MLLMs, which achieve an average of 35.4%. These results highlight a notable difference in how humans and MLLMs process visualizations and are deceived by misleaders. This suggests that the methods needed to make MLLMs more robust to misleading visualizations will, at least in part, differ from those designed to protect human readers (Fan et al., 2022).